OpenStudy (anonymous):

Jernel has to figure out the area of her square garage.she knows that one side of the garage is equal to the length of her rabbit pen. the dimensions of the rectangle rabbit pen are 13 by 10. what is the area of the garage

OpenStudy (anonymous):

Length is going to be the bigger side, so one side of the garage is 13. Since its a square, all sides are equal. Area=l*w Area=13*13

OpenStudy (anonymous):

do u have the answer